### v1

# Conformal Depression Prediction

Abstract
While existing depression recognition methods based on deep learning show promise, their practical application is hindered by the lack of trustworthiness, as these deep models are often deployed as black box models, leaving us uncertain about the confidence of the model predictions. For high-risk clinical applications like depression recognition, uncertainty quantification is essential in decision-making. In this paper, we introduce conformal depression prediction (CDP), a depression recognition method with uncertainty quantification based on conformal prediction (CP), giving valid confidence intervals with theoretical coverage guarantees for the model predictions. CDP is a plug-and-play module that requires neither model retraining nor an assumption about the depression data distribution. As CDP provides only an average performance guarantee across all inputs rather than per-input performance guarantee, we propose CDP-ACC, an improved conformal prediction with approximate conditional coverage. CDP-ACC firstly estimates the prediction distribution through neighborhood relaxation, and then introduces a conformal score function by constructing nested sequences, so as to provide tighter prediction interval for each specific input. We empirically demonstrate the application of uncertainty quantification in depression recognition, and the effectiveness and superiority of CDP and CDP-ACC on the AVEC 2013 and AVEC 2014 datasets.
Index Terms:
depression recognition, uncertainty quantification, conformal prediction, approximate conditional coverage.
I Introduction
Depression is a prevalent mental disorder presented as persistent feelings of sadness, debilitation, and loss of interest in activities [1, 2, 3]. Additionally, depression is often comorbid with many chronic medical diseases, with a substantially higher prevalence among patients affected by diseases such as cardiovascular, metabolic and neurological conditions compared to the general population [4]. It increases the risk of suicide and accounts for a substantial psychological burden [5, 6]. Therefore, AI researchers are working to develop automatic systems for diagnosing depression so that psychiatrists and psychologists can help patients in time. The existing primary diagnostic methods used involve mental health reports, such as the Beck Depression Inventory (BDI-II) [7] (target variable focused on in this paper), the Hamilton Depression Rating Scale (HRSD) [8], and the Patient Health Questionnaire (PHQ-8) [9]. The diagnostic process depends on interviews and demands considerable time and effort from both psychiatrists and patients. It heavily relies on the clinicians’ subjective experience, as well as the patients’ cognitive abilities and psychological states.
Over the past decade, researchers have been exploring effective biomarkers and methods for recognizing depression. In the field of biomarkers, initial attention was focused on hand-crafted features such as Local Phase Quantization (LPQ) [10], Local Gabor Binary Patterns from Three Orthogonal Planes (LGBP-TOP) [11], and Local Binary Pattern from Three Orthogonal Planes (LBP-TOP) [12] extracted from facial videos. Other features explored include speech, facial action units (FAUs), facial landmarks, head poses, and gazes [13, 14]. The development of these hand-crafted features heavily relies on specific knowledge related to the depression recognition task. Acquiring such knowledge is time-consuming, subject to researchers’ subjective cognition and task specificity, lacking good generalization. Moreover, these features struggle to identify certain implicit and difficult-to-distinguish patterns of depression [15]. Fortunately, with the advent of deep learning [16], a new pathway has been opened for depression recognition tasks. Researchers can train end-to-end deep neural networks using depression-related data including facial video [17, 18, 19, 20], speech [21, 22, 13] and other sources. These deep neural networks can discover subtle, indistinguishable depression-related features and implicitly discriminate among them for improved prediction [23].
Figure 1: The pipeline of uncertainty quantification for depression recognition. It can provide a confidence interval for depression prediction at any given confidence level, such as 90%.
While the depression recognition has yielded exciting results, leveraging the rapid development of deep learning and computer vision, its practicality and reliability have consistently raised concerns in clinical applications. Researchers are gradually realizing that deep learning models are black boxes [24, 25]. They lack interpretability, making it difficult to intuitively understand the recognized patterns. Additionally, they lack reliability, as the models may exhibit overconfidence in predicting failure, potentially leading to catastrophic risks. Similarly, depression recognition models also encounter these challenges.
For depression recognition, the predictive performance should not be the only criterion determining the usefulness of a model. The lack of statistically rigorous uncertainty quantification is a key factor undermining the reliability of depression recognition [26]. Consider the application context of depression recognition. We are more willing to accept reliable predictions that meet strict statistical standards and provide insight into potential risks. Therefore, predictive performance should be considered in the context of principled uncertainty quantification. Beyond the depression prediction, we need additional outcomes to represent their uncertainty. It’s imperative to develop uncertainty quantification methods suitable for depression recognition. As shown in Figure 1, our aim is for the model to provide theoretically guaranteed and statistically rigorous confidence intervals in depression predictions. These intervals should cover the actual depression scores with any user-defined level of confidence.
In this work, we propose conformal depression prediction (CDP) to quantify the uncertainty for depression recognition based on conformal prediction (CP [27, 28]). CDP can provide confidence intervals to satisfy marginal coverage for the actual depression scores. In addition, CDP as a plug-and-play module, can validly quantify uncertainty for most depression recognition models, without retraining the model. Furthermore, inspired by adaptive prediction sets (APS) in classification tasks, which can construct a prediction set for input adaptive to its prediction distribution, we propose conformal depression prediction with approximate conditional coverage (CDP-ACC) for providing adaptive confidence intervals for depression recognition to better uncertainty quantification.
The main contributions are summarized as follows:
1. We develop a plug-and-play uncertainty quantification framework CDP for pre-trained depression recognition models. CDP can provide statistically rigorous confidence intervals that satisfy marginal coverage for most existing prediction models. To our knowledge, this is the first attempt to explore uncertainty quantification in depression recognition based on deep neural networks.
2. Furthermore, we propose CDP-ACC, an improved CDP method that satisfies approximate conditional coverage. Different from CDP, CDP-ACC estimates the prediction distribution through neighborhood relaxation, while a new conformal score is applied to ensure tight confidence intervals.
3. We conduct extensive experiments to demonstrate the application of uncertainty quantification in depression recognition and the effectiveness of our proposed method.
II RELATED WORK
II-A Depression Recognition
Over the past decade, the depression recognition models based on deep learning have attracted the interest of many researchers, giving rise to many meaningful researches. Researches focused on depression recognition using labeled depression scores from various modalities including speech, text, video, EEG signals, and the integration of these modalities in a multimodal way have taken a leading role.
For single-modality depression recognition models, He et al. employed a Deep Convolutional Neural Network (DCNN) to extract depression-related features from speech data [21]. These learned features were subsequently integrated with additional hand-crafted features to predict depression scores. Zhao et al. introduced a hybrid network that integrates Self-Attention Networks (SAN) and DCNN to handle both low-level acoustic features and 3D Log-Mel spectrogram data [22]. Following this, feature fusion techniques along with support vector regression are employed for depression prediction. Melo et al. concentrated on efficiently extracting spatiotemporal features from facial videos associated with depression [17, 18]. They introduced a deep Multiscale Spatiotemporal Network (MSN) and a deep Maximization-Differentiation Network (MDN) to predict depression levels in facial videos. Sharma et al. proposed CNN-LSTM hybrid neural networks to extract depression features from EEG signals, which proved to be efficient for depression recognition [29]. Seal et al. developed the DeprNet, a DCNN designed for classification of EEG data from depressed and normal subjects [30].
For multimodal depression recognition models, Niu et al. introduced a novel Spatio-Temporal Attention (STA) network and a Multimodal Attention Feature Fusion (MAFF) strategy to acquire a multimodal representation of depression cues from facial and speech spectrum [31]. Ahmed et al. leveraged features from text, speech, and facial landmarks for multimodal fusion [32]. They proposed an attention-based multimodal classifier with selective dropout and normalization. This approach simplifies the training process on a single neural network. Uddin et al. introduced a novel deep multi-modal framework that effectively integrates facial and verbal cues for automated depression recognition [33]. They extract features separately from video and speech and employ a Multi-modal Factorized Bilinear pooling (MFB) strategy to efficiently fuse these multi-modal features.
These studies primarily concentrate on predicting the severity of depression and emphasize the accuracy of model predictions, while overlooking the analysis of uncertainty. Interestingly, works such as those of Melo et al. introduced minimizing expected error to learn distribution of depression levels [34], while Zhou et al. proposed DJ-LDML to learn depression representation using a label distribution and metric learning [20]. These studies implicitly explore uncertainty in depression recognition, yet without providing a clear and intuitive uncertainty quantification. To our best knowledge, this work is the first attempt to study the uncertainty quantification in depression recognition in modern contexts.
II-B Uncertainty in Affect Computing
Uncertainty has been widely considered for various tasks in affective computing, such as emotion recognition, facial expression recognition, and apparent personality recognition [35, 36, 37, 38, 39, 40, 41]. In emotion recognition tasks, Harper et al. utilized a Bayesian framework to support emotional valence classification from the perspective of uncertainty [35]. Prabhu et al. proposed BNNs to model the label uncertainty based on subjectivity in emotion recognition [36]. Wu et al. employed deep evidence regression to jointly model aleatoric and epistemic uncertainties, aiming to enhance the performance of emotion attribute estimation [37]. Lo et al. utilized probabilistic uncertainty learning to extract information from low-resolution data and proposed the emotion wheel to learn label uncertainty, yielding robust facial expression recognition [38]. In facial expression recognition tasks, She et al. adopted a pairwise uncertainty estimation approach, assigning lower confidence scores to more ambiguous samples to address annotation ambiguities [39]. Le et al. proposed a method based on uncertainty-aware label distribution to adaptively construct the training sample distribution [40]. Additionally, Tellamekala et al. utilized a neural latent variable model to model aleatoric and epistemic uncertainty and integrated both to enhance the performance of apparent personality recognition [41].
The uncertainty-inspired algorithms mentioned above are not suitable for depression recognition, as the available datasets with annotated BDI-II scores are limited. Re-stabilizing a well-performing, uncertainty-inspired depression recognition algorithm is not straightforward. Furthermore, while the above methods may achieve better performance in their respective tasks by introducing uncertainty-inspired heuristics, they do not establish a rigorously effective uncertainty quantification. In contrast to previous methods, our proposed CDP is a plug-and-play uncertainty quantification method for depression recognition, which does not require model retraining and does not assume the depression data distribution. Thanks to the conformal prediction [27, 42], our CDP can achieve valid confidence intervals with theoretical coverage guarantees for model predictions.
However, current conformal prediction methods tend to produce unnecessarily conservative prediction intervals, resulting in the confidence interval to be unnecessarily wide [43]. In depression recognition, overly wide intervals are meaningless. For instance, the prediction for BDI-II (ranging from 0 to 64) has a 100% confidence interval of [0, 64], such a conclusion lacks practicality. Additionally, marginal coverage guarantees only valid average intervals, which overcovers simple subgroups and undercovers difficult ones [44]. Motivated by APS [45], we propose a conformal prediction method with approximate conditional coverage (CDP-ACC) that serves as a plug-and-play solution for the majority of pretrained regression-based depression predictor. Meanwhile, it can produce valid and adaptive confidence intervals based on the specific depression predictions.
III METHOD
III-A Preliminaries
Formally, let us assume that we have the depression-associated data denoted by $X$ as input and the corresponding label denoted by $y$, which is the BDI-II score [7]. We randomly split the data into disjoint training set $\mathcal{D}_{train}$, calibration set $\mathcal{D}_{cal}=\{(X_{i},y_{i})\}_{i=1}^{n}$, and test set $\mathcal{D}_{test}$. We use $\mathcal{D}_{train}$ to train a DNN model $\hat{f}(\cdot)$ for depression recognition. Here, $\hat{y}\in\mathbb{R}$ denotes the prediction for $X$ by depression recognition model $\hat{f}(\cdot)$. We can formulate depression recognition as a vanilla regression task by minimizing the MSE loss (Eq. 1) as the learning objective.
$\begin{split}\mathcal{L}_{MSE}=\frac{1}{2N}\sum_{i=1}^{N}(\hat{y}_{i}-y_{i})^{2}\end{split}$ (1)
In depression recognition based on vanilla regression models, the depression prediction $\hat{y}$ is the expectation $\mathbb{E}[\mathcal{P}(y|X)]$ derived from estimating the conditional probability distribution $\mathcal{P}(y|X)$. However, We do not have access to the oracle of $\mathcal{P}(y|X)$. To quantify the uncertainty of depression prediction $\hat{y}$, we expect to construct an interval function $\mathcal{C}_{\alpha}(\cdot)$, which takes depression data and miscoverage rate $\alpha$ as input and outputs a confidence interval. The interval contains the depression target $y$ with any confidence level $1-\alpha$, as shown in Fig 1 ($\alpha=0.1$). Given the coverage of $1-\alpha$, $|\mathcal{C}(X_{n+1})|$ represents the width of the interval $X_{n+1}$, which indicates the uncertainty of $\hat{y}_{n+1}$. A wider interval $|\mathcal{C}(X_{n+1})|$ indicates greater uncertainty in the model’s depression prediction $\hat{y}_{n+1}$.
Figure 2: The pipeline of CDP. CDP requires a pre-trained depression recoginition model $\hat{f}$, a calibration set $\mathcal{D}_{cal}$ and a confidence level $1-\alpha$. As a post-processing method for uncertainty quantification, we utilize the depression predictions of the calibration set to calculate conformal scores. Subsequently, we employ the $(1-\alpha)$-th empirical quantile value $\hat{s}$ of conformal scores to form confidence intervals for $X_{test}$.
III-B Depression Recognition with UQ
The straightforward method for uncertainty quantification in depression recognition is to utilize a DNN to estimate the conditional probability distribution $\mathcal{P}(y|X=x)$ of depressed individuals. Given our lack of knowledge regarding the true $\mathcal{P}(y|X=x)$, currently, the most common approach assumes that depression target $y$ follow a Gaussian distribution $y\sim\mathcal{G}(\hat{y},\sigma^{2})$. Using the negative log-likelihood (NLL) for the Gaussian distribution [46], we can derive the loss function:
$\mathcal{L}_{NLL}=\sum_{i=1}^{N}\frac{log\sigma_{i}^{2}}{2}+\frac{(y_{i}-\hat{y}_{i})^{2}}{2\sigma_{i}^{2}}$ (2)
Minimizing Eq 2 as the learning objective during training allows us to obtain a DNN capable of predicting the distribution $\mathcal{P}(y|X=x)$ for any given individual. Furthermore, in the context of the Gaussian distribution, the quantile can be calculated as follows:
$\hat{y}_{\alpha}=\hat{y}+\sigma\cdot\mathrm{erf}^{-1}(\alpha)$ (3)
where $\mathrm{erf}^{-1}$ denotes inverse Gaussian error function. we can acquire the confidence interval $[\hat{y}_{\alpha/2},\hat{y}_{1-\alpha/2}]$, that satisfies $\mathcal{P}(y\in[\hat{y}_{\alpha/2},\hat{y}_{1-\alpha/2}])\geq 1-\alpha$.
In practice, however, depression datasets [47, 48] exhibit high data imbalance, and the assumption of the Gaussian distribution may not be appropriate in this context. We prefer a distribution-free method for uncertainty quantification. Fortunately, with the introduction of pinball loss (Eq 4), we can train a DNN as a conditional quantile estimator [49, 50], allowing us to obtain arbitrary quantile values for $\mathcal{P}(y|X=x)$. Then, we can use ${\alpha/2}$ and ${1-\alpha/2}$ quantile values to derive the confidence interval $[\hat{y}_{\alpha/2},\hat{y}_{1-\alpha/2}]$, thereby achieving a distribution-free uncertainty quantification method.
Input: Calibration set $\mathcal{D}_{cal}=\{X_{i},y_{i}\}_{i=1}^{n}$, pre-trained depression recoginition model $\hat{f}(\cdot)$, test data $X_{n+1}$, miscoverage rate $\alpha\in(0,1)$, quantile estimator $\mathcal{Q}(\cdot,\cdot)$
Output: Confidence interval $\mathcal{C}_{\alpha}(X_{n+1})$
1
// Calculate the conformal scores $\{s_{i}\}_{i=1}^{n}$ on $\mathcal{D}_{cal}$
2 for $i=1$ to $n$ do
3      $\hat{y}_{i}=\hat{f}(X_{i})$
4       $s_{i}=|\hat{y}_{i}-y_{i}|$
5 end for
6
[…]
// Calculate the $q$-th quantile of the conformal scores
8
[…]
// Calculate the confidence interval
10 $\mathcal{C}_{\alpha}(X_{n+1})=[\hat{f}(X_{n+1})-\hat{s},\hat{f}(X_{n+1})+\hat{s}]$
return $\mathcal{C}_{\alpha}(X_{n+1})$
Algorithm 1 Conformal Depression Prediction
$\mathcal{L}_{Pinball}=\begin{cases}q\cdot(y-\hat{y}_{q}),&\text{if }y\geq\hat{y}_{q}\\ (1-q)\cdot(\hat{y}_{q}-y),&\text{if }y<\hat{y}_{q}\end{cases}$ (4)
The prerequisite for implementing the above method is that we can obtain a well-trained DNN model. However, due to limited depression datasets and imbalanced depression data distributions, using Eq 2 or Eq 4 to train depression recognition models is challenging. In practice, it has been observed that those methods suffer from overconfidence, where $\mathcal{P}(Y_{i}\in C(X_{i}))$ is significantly lower than $1-\alpha$. In this context, there is an urgent need for a post-processing method that is distribution-free to quantify the uncertainty of depression recognition. In the application of depression recognition, it’s impossible to access the conditional distribution $\mathcal{P}(y|X)$. To address this challenge, we can ensure that $\mathcal{C}(\cdot)$ guarantees marginal coverage to maintain an average coverage rate across the target, as expressed in Eq 5.
$\mathcal{P}(y_{test}\in\mathcal{C}(X_{test}))\geq 1-\alpha,\ \ \ (X_{test},y_{test})\in\mathcal{D}_{test}$ (5)
Conformal prediction (CP [27, 51]) addresses the requirement for uncertainty quantification in the field of depression recognition. We develop conformal depression prediction (CDP), which applies CP to the general depression recognition algorithm, as shown in Figure 2. The specific steps of CDP are shown in Algorithm 1. With CDP, we can acquire valid, theoretical coverage guaranteed confidence interval, and adjust the interval width according to user-defined confidence levels. We can ensure that this confidence interval achieves marginal coverage for the target depression scores at any confidence levels, as proven by [27, 51].
III-C CDP with Approximate Conditional Coverage
There is a stronger notion 6 than the marginal coverage, namely conditional coverage.
$\mathcal{P}(y_{n+1}\in\mathcal{C}(X_{n+1})|X_{n+1})\geq 1-\alpha$ (6)
for every input $X_{n+1}$, we aim to provide a confidence interval with $1-\alpha$ coverage. In contrast to marginal coverage, conditional coverage is more suitable for uncertainty quantification in individual depression prediction. In depression recognition applications, each individual only cares about whether the recognition prediction is accurate (or within which confidence interval does the depression score fall). Unfortunately, achieving conditional coverage is practically challenging in the finite datasets and a distribution-free setting [42]. Our work draws inspiration from ordinal adaptive prediction sets [26] in classification tasks. The motivation behind CDP-ACC is to generate adaptive confidence intervals that achieve approximate conditional coverage (Eq 6) in a model-agnostic and distribution-free manner for pretrained depression predictor.
Figure 3: The pipeline of CDP-ACC. Following the steps of CDP, CDP-ACC also requires a calibration set. We partition the predictions of the calibration set into bins. For samples within each bin, we estimate the distribution of depression scores $y$ using histograms, and calculate the conformal scores within the group according to Eq 7.
Approximate Conditional Coverage. In pretrained depression predictor, $\hat{y}=\mathbb{E}[\mathcal{P}(y|X)]$, we cannot directly access to the distribution $\mathcal{P}(y|X)$. To accurately estimate $\mathcal{P}(y|X)$ in depression recognition, we rely on a key assumption from [52]: similar $X$ will have similar conditional distributions $\mathcal{P}(y|X)$. Formally, if $X_{i}\approx X_{j}$, $\mathcal{P}(y|X=X_{i})\approx\mathcal{P}(y|X=X_{j})$, that is $\mathbb{E}[\mathcal{P}(y|X_{i})]\approx\mathbb{E}[\mathcal{P}(y|X_{j})]$. This assumption is also reasonable in our practical observations. For example, different video clips of the same patient tend to have similar depression predictions.
Then, suppose there are samples $X$ belonging to a very small neighborhood $\Delta(X_{i})$, which we can consider they share the similar conditional distribution according to the smooth assumption above. The straightforward idea to approximate $\mathcal{P}(y|X=X_{i})$ is to replace $X=X_{i}$ with $X\in\Delta(X_{i})$, that is $\mathcal{P}(y|X=X_{i})\approx\mathcal{P}(y|X\in\Delta(X_{i}))$. In the context of facial depression recognition, $X_{i}$ denotes facial video clips. $\Delta(X_{i})$ can represent clips belonging to the same video, or clips with the same depression score. However, for unseen test video clips, we have no oracle with access to $\Delta(X_{i})$. We need to further explore relaxation to approximate $\Delta(X_{i})$.
We revisit $\Delta(X_{i})$ from the perspective of the model (pre-trained model $\hat{f}$): $\Delta(\hat{y_{i}})=\hat{f}(\Delta(X_{i}))$. The model predicts similar depression scores for what it perceives as similar $X_{i}$. Thus, we can use $\Delta(\hat{y_{i}})$ as the condition instead of $\Delta(X_{i})$. That is, we further assume based on [52]: similar $\hat{y}$ also have similar conditional distributions [^1] This assumption can be understood as follows: we all know that the mean squared error (MSE) loss can be derived as the maximum likelihood estimate of a Gaussian distribution with a mean of $\hat{y}$ and a fixed variance. Therefore, if $\hat{y}$ is similar, then the fixed-variance Gaussian distribution will be similar.. Then, $\mathcal{P}(y|X=X_{i})\approx\mathcal{P}(y|\hat{y}\in\Delta(\hat{y}))$. Given the smoothness assumption (demo shown in Figure 4), we can estimate the conditional distribution by grouping neighboring points of the depression predictions.
Figure 4: Demo of smoothed assumption. We measure the uncertainty of the model predictions by considering $\mathcal{P}(y|\hat{y}\in\Delta(\hat{y})$. For example, we consider all individuals with depression predictions $\hat{y}=1\sim 3$ as a group. If the model’s predictions are reliable, the actual depression scores for this group would be about $1\sim 3$. If the model’s predictions are unreliable, then there will be abnormal depression scores in this group.
Conformal score. In conformal prediction, the setting of conformal score is particularly important. Different conformal scores essentially determine the effectiveness of the conformal method. In CDP-ACC, the setting of conformal score consists of dividing bins and constructing a nested sequence of approximate oracle intervals.
In $\mathcal{D}_{cal}$, we first divide $\hat{y}\in[0,K)$ into $M$ bins: $B_{m}=[\frac{m-1}{M}K,\frac{m}{M}K)$, where $m=1,...,M$ and $K$ denotes the predetermined upper bound. Now, we have $M$ bins and predictions $\hat{y}\in B_{m}$ fall in bins. Based on the approximation $\mathcal{P}(y|X=X_{i})\approx\mathcal{P}(y|\hat{y}\in B_{m})$, we only require the depression label $y$ corresponding to $\hat{y}$ within $B_{m}$ to approximately obtain the conditional distribution $\mathcal{P}(y|X=X_{i})$.
Subsequently, we need to estimate the distribution of $y$ within each bin. Referring to [51, 26], we use histogram estimation to acquire $\mathcal{P}_{m}(y|\hat{y}\in B_{m})$ for bin $m$. A good confidence interval should achieve the specified coverage rate. Narrower interval widths are desirable as they provide more precise uncertainty quantification.
To maintain the continuity of prediction interval, we construct a nested sequence $E(\cdot)$ so that the interval with higher confidence covers the interval with lower confidence, all while endeavoring to minimize the width of each interval, as shown in Figure 5.
$E_{m}(\tau,\mathcal{P}_{m})=\operatorname*{argmin}_{[y_{lo},y_{hi}]\subseteq{B_{m}}}\{y_{hi}-y_{lo}:\mathcal{P}_{m}(y\in[y_{lo},y_{hi}])\geq\tau\}$ (7)
where $\tau\in(0,1)$ and $m\in[1,M]$. If the optimal solution is not unique, we can add a little noise into the $B_{m}$ to break the ties. It should note that $E_{m}(\cdot)$ is a nested sequence for bin $m$, for example $E_{m}(0.1)\subseteq E_{m}(0.2)\subseteq E_{m}(0.3)$, and $E_{m}(\tau)$ is a interval $[y_{m,lo},y_{m,hi}]$.
Figure 5: The specific process of conformal score. First, bins are divided according to $\hat{y}$, and we use each bin to denote the neighborhood of $\hat{y}$. Second, we compute the sub-conformal-score by histogram estimation within each bin. Finally, the predicted interval can be acquired based on the bin assigned to the test sample $\hat{f}(X_{n+1})$.
Confidence Interval. After calculating the conformal scores in the calibration set, we can leverage the assumption that the calibration set and the test samples are exchangeable to generate valid confidence intervals. Specifically, given all conformal scores $E_{m}$ in the calibration set, a user-defined miscoverage rate $\alpha$, and an unseen test sample $X_{n+1}$, CDP-ACC acquires confidence intervals according to the following steps. First, for each bin $m$, calculate the $(1-\alpha)$th quantile of the conformal scores using Eq. 7, and acquire the corresponding interval $[y_{m,lo},y_{m,hi}]$. Within each bin of the calibration set, there exists a confidence interval with a coverage rate of $1-\alpha$ and the shortest width. Then, use the pre-trained model $\hat{f}$ to predict the depression score of $X_{n+1}$, i.e., $\hat{y}=\hat{f}(X_{n+1})$. The corresponding confidence interval $[y_{m,lo},y_{m,hi}]$ can be obtained based on the bin $m$ into which $\hat{y}_{n+1}$ falls.
Input: Calibration set $\mathcal{D}_{cal}=\{X_{i},y_{i}\}_{i=1}^{n}$, pre-trained depression recoginition model $\hat{f}(\cdot)$, test data $X_{n+1}$, miscoverage rate $\alpha\in(0,1)$, upper bound $K$, histogram estimator $\mathcal{H}(\cdot)$, bins number $M$
1
Output: Confidence interval $\mathcal{C}_{\alpha}(X_{n+1})$
2 Initialize bins $\{B_{m}\}_{m=0}^{M-1}\leftarrow\emptyset$
// Bins dividing for $\hat{y}$ on $\mathcal{D}_{cal}$
3
4for $i=1$ to $n$ do
5       $\hat{y}_{i}=\hat{f}(X_{i})$
6      for $m=1$ to $M$ do
7             if $\hat{y}_{i}\in[\frac{m-1}{M}K,\frac{m}{M}K)$ then
8                   $B_{m}\cup\{y_{i}\}$
9             end if
10
11       end for
12
13 end for
// Histogram estimation
14 for $m=1$ to $M$ do
15       $\mathcal{P}_{m}(y|\hat{y}\in B_{m})$ = $\mathcal{H}(B_{m})$
16 end for
17
// Calculate the confidence interval
18
[…]
21 for $m=1$ to $M$ do
       // Calculate subinterval according to Eq 7
22       $[y_{m,lo},y_{m,hi}]=E_{m}(q,\mathcal{P}_{m})$
23       if $\hat{y}_{n+1}\in B_{m}$ then
24             $\mathcal{C}_{\alpha}(X_{n+1})=[y_{m,lo},y_{m,hi}]$
25       end if
26
27 end for
28
29return $\mathcal{C}_{\alpha}(X_{n+1})$
Algorithm 2 Depression Recognition with CDP-ACC
IV EXPERIMENTS
IV-A Data
In order to explore uncertainty quantification methods in depression recognition, we work on two classic facial depression datasets, AVEC 2013 [47] and AVEC 2014 [48]. AVEC 2013 is a subset of the AViD-Corpus, consisting of 150 facial videos. The videos range in length from 20 to 50 minutes, with a frame rate of 30 FPS and a resolution of 640 * 480 pixels. The content of the videos includes facial recordings of subjects during human-computer interaction tasks, such as reading assigned content and improvising on given sentences. We divide the dataset into 70 videos for training, 30 videos for validation (calibration), and 50 videos for testing. The facial depression levels are annotated using BDI-II scores [7], with each video assigned a single value representing the depression level. The BDI-II questionnaire consists of 21 questions, each with multiple-choice answers scored from 0 to 3. The total BDI-II score ranges from 0 to 63: 0-13 indicates minimal depression, 14-19 indicates mild depression, 20-28 indicates moderate depression, and 29-63 indicates severe depression.
AVEC 2014, a subset of AVEC 2013, consists of two tasks: Northwind and Freeform. It comprises 150 Northwind videos and 150 Freeform videos, totaling 300 videos. The original videos of AVEC 2014 are similar to those used in the AVEC 2013 but include five pairs of previously unseen videos, replacing a few videos considered unsuitable. We use 200 videos for training, 50 videos for validation (calibration set), and 50 videos for testing.
It is ensured that the three datasets had similar distributions in terms of age, gender, and depression severity. There is no overlap between the three datasets. Depression severity labels are still based on BDI-II scores [7]. The calibration set should meet the exchangeability with the test set [53]. The better the calibration set, the better the performance of CDP. The videos in the AVEC 2014 dataset are relatively short, with an overlap of 8 frames between adjacent video clips, whereas there is no overlap in the AVEC 2013 dataset. During data processing, landmarks in all video frames are detected using OpenFace [54], and facial region alignment is performed using the eyes, nose, and mouth.
IV-B Experimental Setup
Predictive Model. We focus on post-processing methods and do not have strict requirements on the model. For the purpose of our research without loss of generality, we adopt the classic C3D [55, 19] network as the predictive model for depression recognition. A slight adjustment to the fully connected (FC) layer could better suit the task [20]. In particular, for vanilla regression, we employ MSE Loss as the learning objective during training and modify the last FC layer from 4096 to 64 to enhance training stability and prevent overfitting. In the context of Quantile Regression (QR [49]), we use 99 quantiles evenly spaced between 0.01 and 0.99 for the pinball loss. Then, the output of the last FC layer transforms from a single output in vanilla regression to 99 outputs in QR. To validate the effectiveness of CDP-ACC in the field of depression recognition, we also employ SlowFast [56] as a backbone.
UQ for Depression Recognition. To quantify the uncertainty of depression recognition models, we perform NLL [46] and QR [49] that achieves UQ by retraining the model. We also implement several post-processing methods related to CP [27, 42] to construct valid confidence intervals for depression recognition. Specifically, CQR [43] is a post-processing method that applies conformal prediction to calibrate the interval of pre-trained QR. Its conformal score is $max\{\hat{f}(X;\alpha/2)-y,y-\hat{f}(X;1-\alpha/2)\}$. CDP represents the method for uncertainty quantification in vanilla regression, as shown in Figure 2. Its conformal score is defined as $|y-\hat{y}|$, which achieves marginal coverage. CDP-ACC is the method that achieves approximate conditional coverage and provides valid and adaptive confidence intervals. We set the miscoverage rate $\alpha$ to 0.1. Given that the depression score range is from 0 to 63, we use truncation to limit $\hat{y}_{lo}$ to not be less than 0, and $\hat{y}_{hi}$ not to be greater than 63. For CDP-ACC, we set bins number $M$ to 10 and upper bound $K$ to 63. In addition, we also conduct UQ experiments on seven publicly available general regression datasets (for more information, please refer to the supplementary materials).
Metrics. In depression recognition, we commonly use mean absolute error (MAE) and root mean squared error (RMSE) as evaluate metrics for predicting errors, aiming to quantify the accuracy of the prediction results. To further measure the inherent uncertainty in predictions, we apply the prediction interval coverage probability (PICP) and mean prediction interval width (MPIW) to assess the uncertainty of confidence intervals. PICP and MPIW are two seemingly contradictory metrics. For example, if the PICP worsens, the MPIW may actually improve. We aim to minimize MPIW while ensuring PICP coverage rate. However, if we cannot ensure that PICP is not less than (1 - $\alpha$) $\times$ 100%, merely shortening the interval width is meaningless. Similarly, constructing an interval with a larger width to ensure coverage is also meaningless. For instance, we all know that at 100% confidence level, the interval for the patient’s depression prediction is [0, 64], which is correct but meaningless. With the condition that PICP satisfies $1-\alpha$, a smaller MPIW is considered more desirable.
$PICP=\frac{1}{|\mathcal{D}_{test}|}\sum_{i}^{|\mathcal{D}_{test}|}\mathbf{1}(y_{i}\in\mathcal{C}(X_{i}))\times 100\%,\ \ X\in|\mathcal{D}_{test}|$ (8)
$MPIW=\frac{1}{|\mathcal{D}_{test}|}\sum_{i}^{|\mathcal{D}_{test}|}|\hat{y}_{i,hi}-\hat{y}_{i,lo}|,\ \mathcal{C}(X_{i})=[\hat{y}_{i,lo},\hat{y}_{i,hi}]$ (9)
TABLE I: Results of prediction error
Method Backbone Dataset MAE RMSE
Vanilla C3D AVEC 2013 7.14 8.97
NLL [46] C3D AVEC 2013 7.02 8.92
QR [49] C3D AVEC 2013 6.82 8.64
Vanilla C3D AVEC 2014 6.54 8.32
NLL [46] C3D AVEC 2014 7.10 9.24
QR [49] C3D AVEC 2014 6.65 8.24
Vanilla SlowFast AVEC 2013 7.49 9.37
NLL [46] SlowFast AVEC 2013 7.30 9.21
QR [49] SlowFast AVEC 2013 7.22 9.04
Vanilla SlowFast AVEC 2014 7.00 8.70
NLL [46] SlowFast AVEC 2014 6.29 8.11
QR [49] SlowFast AVEC 2014 6.89 8.65
TABLE II: Results of uncertainty quantification
Method Backbone Dataset PICP MPIW
NLL [46] C3D AVEC 2013 74.33% 13.17
QR [49] C3D AVEC 2013 17.61% 4.73
CQR [43] C3D AVEC 2013 87.94% 23.32
CDP C3D AVEC 2013 91.78% 22.82
CDP-ACC C3D AVEC 2013 92.00% 21.48
NLL [46] C3D AVEC 2014 76.52% 14.15
QR [49] C3D AVEC 2014 21.84% 5.35
CQR [43] C3D AVEC 2014 87.14% 23.53
CDP C3D AVEC 2014 91.53% 27.47
CDP-ACC C3D AVEC 2014 91.67% 22.75
NLL [46] SlowFast AVEC 2013 66.94% 17.69
QR [49] SlowFast AVEC 2013 22.30% 5.38
CQR [43] SlowFast AVEC 2013 92.29% 27.78
CDP SlowFast AVEC 2013 95.25% 28.29
CDP-ACC SlowFast AVEC 2013 91.25% 24.13
NLL [46] SlowFast AVEC 2014 69.57% 19.59
QR [49] SlowFast AVEC 2014 70.49% 16.83
CQR [43] SlowFast AVEC 2014 93.06% 27.76
CDP SlowFast AVEC 2014 93.86% 26.28
CDP-ACC SlowFast AVEC 2014 92.18% 23.29
$\alpha=0.1$
Furthermore, we utilize size-stratified coverage (SSC [57, 44]) to assess the extent to which different methods achieve conditional coverage for subsets of varying depression severity. We divide all possible test data into $G$ bins based on the severity of depression (shown in Table IV), denoted as $B_{1},B_{2},\ldots,B_{G}$. The formula for SSC is as follow:
$SSC=\underset{g\in\{1,..,G\}}{\min}\frac{1}{|\mathcal{I}_{g}|}\sum_{i\in\mathcal{I}_{g}}{\mathbf{1}\{y_{i}\in\mathcal{C}(X_{i})\}}$ (10)
where $\mathcal{I}_{g}\subset\{1,2,\ldots,n_{\text{test}}\}$ represents the observations falling into bin $g$, $|\mathcal{I}_{g}|$ denotes the size of $\mathcal{I}_{g}$ and $\mathbf{1}(\cdot)$ denotes indicator function.
IV-C Results
In this section, we provide a comprehensive evaluation of uncertainty quantification methods for depression recognition through experimental results and analysis. Firstly, Table I displays the prediction errors for vanilla regression, NLL and QR, allowing for a comparison of their accuracy in depression recognition. Furthermore, we compare the uncertainty quantification performance of different UQ methods in depression recognition (shown in Table II), demonstrating the superiority of CDP-ACC. We present in Table III the approximation level of different methods for conditional coverage. We also demonstrate the uncertainty analysis of CDP-ACC for some samples with considerable errors to further enhance insight into CDP-ACC (shown in Table V). Additionally, in the supplementary material, we plot confidence intervals for different methods, providing an intuitive way to discern the strengths and weaknesses of each method.
TABLE III: Results of conditional coverage on AVEC 2014
Method Backbone SSC
NLL [46] C3D 0.2426
QR [49] C3D 0.1715
CQR [43] C3D 0.8280
CDP [27] C3D 0.7035
CDP-ACC C3D 0.8980
NLL [46] SlowFast 0.5293
QR [49] SlowFast 0.4623
CQR [43] SlowFast 0.8343
CDP [27] SlowFast 0.8140
CDP-ACC SlowFast 0.8407
$\alpha=0.1$
Table I shows the prediction errors of the C3D and SlowFast under different datasets and methods. We can find that comparing vanilla regression using MSE or NLL as the learning objective. The prediction error for quantile regression is slightly smaller in AVEC 2013. This indicates the 0.5 quantile value of QR has a smaller error than the vanilla regression and NLL. However, in the experiments with the SlowFast model in AVEC 2014, the NLL yields smaller errors compared to the other two methods. This also indicates that if only prediction errors are considered, the NLL method as well as the QR may present better performance. However, both methods demonstrate instability in training processes and results in depression recognition.
Table II shows the application of several uncertainty quantification methods in different depression recognition models. Due to NLL and QR heavily relying on the training process, it’s intuitive to observe that the coverage of QR is the worst, followed by NLL. In the current context of depression recognition, inevitable overfitting, unstable training process and imbalanced datasets make it difficult to obtain effective quantile values or fitting Gaussian distributions. For pre-trained QR, CQR is a valid method to calibrate coverage by adjusting the interval width. In this way, CQR can provide adaptive confidence intervals, but its adaptiveness primarily stems from QR, while the conformal prediction mainly provides theoretical guarantees for marginal coverage. In situations where QR exhibits overconfidence, CQR unavoidably relies on average marginal coverage. As for CDP-ACC, it is an approximate condition coverage method for vanilla regression. It can reduce the interval width as much as possible by providing adaptive confidence intervals while maintaining coverage.
Table III displays the conditional coverage levels of different UQ methods (with $\alpha=0.1$) for various levels of depression severity, which are categorized into four groups (shown in Table IV). Therefore, a SSC closer to $1-\alpha$ indicates better conditional coverage of the UQ method for different levels of depression severity. We can clearly observe that NLL and QR perform the worst in terms of SSC due to unstable training and overconfidence. As an effective method for quantifying depression uncertainty, CDP can only guarantee marginal coverage, hence its results are not as good as CQR. However, the conditional coverage of CQR is limited by the pre-trained QR model. In contrast, CDP-ACC can directly approximate conditional coverage based on the predictions of vanilla regression models, which does not require retraining the QR model and achieves better conditional coverage.
TABLE IV: The Relation Between BDI-II and Depression Severity Level
BDI-II Score Severity Level
0-13 minimal
14-19 mild
20-28 moderate
29-63 severe
TABLE V: depression recognition with uncertainty quantification
$\mathbf{Label\ y}$ 11 11 33 43
$\mathbf{C3D\ \hat{y}}$ 0.56 18.4 19.9 25.1
$\mathbf{Interval}$ [0,11] [11,33] [11,33] [10,43]
In the Table V, we present several specific examples with notable errors in depression predictions. Solely relying on the depression prediction of the model, we cannot acquire the model’s confidence in the prediction (e.g., whether the model correctly recognized depression). CDP-ACC, a post-processing method, provides valid confidence intervals for the model’s predictions. A wider interval indicates potentially lower confidence in the model’s predictions. In addition, it also provides us with insights into assessing potential risks of depression.
In the context of depression recognition, we require theoretically guaranteed, statistically rigorous uncertainty quantification methods. Through experimentation, we have observed that methods such as retraining models using NLL and QR tend to exhibit overconfidence. Therefore, we expect post-processing methods and propose CDP for uncertainty quantification by providing confidence intervals. Additionally, we have noticed that CQR can calibrate pre-trained QR and provide adaptive prediction intervals. However, the majority of depression recognition methods are currently based on vanilla regression. Inspired by adaptive prediction sets in classification, we introduce CDP-ACC. By achieving approximate conditional coverage, CDP-ACC offers a plug-and-play uncertainty quantification method for pretrained regression-based depression predictor.
V CONCLUSION
The lack of statistically rigorous uncertainty quantification methods results in unreliable depression recognition. By introducing conformal prediction, we develop a plug-and-play uncertainty quantification method CDP to produce confidence intervals for depression recognition. The intervals predicted by CDP can achieve marginal coverage guarantee at any miscoverage rate. Furthermore, akin to the adaptive prediction sets in classification tasks, we introduce CDP-ACC for depression recognition. It alleviates the issue of marginal coverage practices focusing solely on the average interval width while overlooking individuals with great uncertainty by achieving approximate conditional coverage, and provides valid and adaptive confidence intervals for better uncertainty quantification in depression recognition. We expect that this work will attract more researchers to pay attention to the issue of uncertainty quantification in depression recognition.
V-A Limitations
CDP still faces some challenges in practice:
1. Exchangeable calibration set: The effectiveness of CDP relies entirely on the assumption that the calibration set and test samples are exchangeable, swapping test samples with samples from a calibration set does not change the distribution of data in the calibration set. In practical usage, while we can use a validation set that is independent and identically distributed (IID) with the test set as the calibration set, achieving IID in many cases is challenging. Once there is a distribution shift in the test set, the effectiveness of CDP may diminish.
2. Interval width remains wider than expected: CDP can only provide valid intervals as a post-processing method and does not impact the training process. Currently, the error in depression recognition tasks remains considerable, resulting in wider intervals, indicating less confidence in the model’s results.
V-B Ethical Statement
As a post-processing method, CDP does not compromise user privacy. It is important to note that uncertainty quantification in depression recognition needs to be conducted with user consent, and there should be a thorough explanation of the potential for model prediction failures and the possibility of underlying depression risks.
Acknowledgments
This work was funded in part by the National Natural Science Foundation of China under Grants 61972046.